OpenStudy (anonymous):

The function below does not have an inverse. Restrict the domain of the function so that the restricted function has an inverse. f(x) = 3(x + 5)2 - 1 Restricted domain. x ≥

OpenStudy (anonymous):

What am I supposed to do here?

OpenStudy (anonymous):

That is a "squared," yes?

OpenStudy (anonymous):

o yes. it is. my bad.

OpenStudy (anonymous):

f(x)=3(x+5)^2-1

OpenStudy (anonymous):

You need to find the vertex of that parabola and only use one side of it or the other... When you invert it, you end up with more than one y value for some of the x values.

OpenStudy (anonymous):

x>=-5

OpenStudy (anonymous):

so use, x= -b/(2a) to find the x-value of the vertex then you can use the graph above that point.

OpenStudy (anonymous):

o and the last part should be x ≥ ___ (blank)

OpenStudy (anonymous):

oh - oops... it's already in the right format LOL yes, x=-5

OpenStudy (anonymous):

could you explain/show work please

OpenStudy (anonymous):

so half of the parabola is to the right of that x-value and the other half is to the left. Choose the part you are asked for and when you invert it, there will be no problem with it being a function

OpenStudy (anonymous):

f(x)=3(x+5)^2-1 the vertex is (-5,-1) because it is in the format y=n(x-h)^2+k where (h,k) is the vertex

OpenStudy (anonymous):

i calculated the inverse . its has sqrt(y+1/3) which should be greateer than 0 for gods sake no negative numbers under root. so if y>-1 corresponding expression also greater than -1. so solve for x

OpenStudy (anonymous):

the number is positive 5... how did get a negative 5?

OpenStudy (anonymous):

@mrslaybms

OpenStudy (asnaseer):

@mariomintchev do you understand why the function given, i.e.: \[f(x)=3(x+5)^2-1\]does NOT have an inverse?

OpenStudy (anonymous):

because its squared?

OpenStudy (asnaseer):

but do you what that "means"?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

have you talked about the 'vertical line test' in class at all?

OpenStudy (anonymous):

o yeah

OpenStudy (anonymous):

cant touch more than one point when a vertical line is drawn

OpenStudy (anonymous):

in response to the 5 vs -5 question, Mario... the formula is (x-h) so what number is subtracted from the x in (x-5)? It's just 5, not -5.

OpenStudy (asnaseer):

1. the function will return some value for every value of x you put into it. 2. however, there are pairs of values for x which will give you the same value for the function (this is because the function contains a squared term as you suggested) 3. this means that if you know the value of the function, there is no way of knowing which of the pair of x values it came from this is what makes it not have an inverse

OpenStudy (anonymous):

so the answer was -5 because its the inverse?

OpenStudy (anonymous):

in response to the vertical line test, exactly! :) SO,,,, if the original graph, before you find it's inverse has more then one point on a HORIZONTAL line, it's inverse will have more than one point on a VERTICAL line. Does that make sense?

OpenStudy (anonymous):

o cause it can be 2 positives that give u the same answer and 2 negatives that can give u the same answer...

OpenStudy (anonymous):

2 * 2 = 4 and -2 * -2 = 4

OpenStudy (anonymous):

@mrslaybms i guess

OpenStudy (anonymous):

Here are a couple other examples to help you with the 5 part... If y =(x+3)^2-1, it's vertex would be (-3,-1). You use the opposite of the number in the ( ) and the exact number on the outside. So... if y=(x-6)^2+3, it's vertex would be (6,3) and y=(x-2)^2-4 would be (2,-4).

OpenStudy (anonymous):

2 points 2 answers... which means its not a one-on-one function... im guessing...

OpenStudy (anonymous):

i see i see

OpenStudy (asnaseer):